# Optimal network sizes for most robust Turing patterns

### v1

@@ Abstract
Many cellular patterns exhibit a reaction-diffusion component, suggesting that Turing instability may contribute to pattern formation. However, biological gene-regulatory pathways are more complex than simple Turing activator-inhibitor models and generally do not require fine-tuning of parameters as dictated by the Turing conditions. To address these issues, we employ random matrix theory to analyze the Jacobian matrices of larger networks with robust statistical properties. Our analysis reveals that Turing patterns are more likely to occur by chance than previously thought and that the most robust Turing networks have an optimal size, surprisingly consisting only of a handful of molecular species, thus significantly increasing their identifiability in biological systems. This optimal size emerges from a tradeoff between the highest stability in small networks and the greatest instability with diffusion in large networks. Furthermore, we find that with multiple immobile nodes, differential diffusion ceases to be important for Turing patterns. Our findings may inform future synthetic biology approaches and provide insights into bridging the gap to complex developmental pathways.
@@ Introduction
Spatial patterns and structures are ubiquitous in biological systems, ranging from microbial communities and biofilms to developmental biology and ecological systems [1] [2] [3]. In seminal work, Alan Turing (and later Gierer-Meinhardt) proposed a self-organizing emergent mechanism for pattern formation based on a diffusion-driven instability, utilizing a slowly diffusing activator and a fast-diffusing inhibitor molecule [4] [5]. According to Turing’s definition, the chemical reactions are stable without diffusion, leading to a homogeneous steady state, but become unstable with diffusion, forming a periodic pattern at certain wave numbers. The mathematical analysis is generally based on reaction-diffusion models, implemented using partial differential equations, and linear stability analysis to investigate the effect of perturbations on a steady state [6]. Nonetheless, employing Turing models to understand biological patterns is still a topic of debate.
@@
The Turing mechanism clearly has biological relevance in cell and developmental biology, likely explaining aspects of digit formation in mice, zebrafish skin, fingerprint formation, and cortical folds of the fetal brain [7][8][9][10]. However, this mechanism has two main drawbacks: extreme simplicity and a need for fine-tuning. In contrast to simple activator-inhibitor Turing models, biological gene regulatory networks in embryonic development generally consist of hundreds of molecular species and are notable for their immense complexity, hierarchical structure, and tolerance to noise [11]. In contrast, the Turing conditions lead to a lack of structural robustness, which is at odds with the noise tolerance and evolutionary adaptability required for such patterning solutions to occur [12][13][14]. However, there are also bottom-up approaches to further elucidate the issues with Turing models.
@@
As developmental biology is complex, synthetic biology provides an alternative route, following Feynman’s mantra: “What I cannot create, I do not understand” [15]. By implementing the Turing mechanism from scratch in cells that communicate via chemicals, our understanding of how to generate stable pattern can be systematically improved. Two previous synthetic biology attempts failed due to the lack of differential diffusion, leading to highly irregular patterns. First, stochastic Turing patterns were engineered in E. coli cells. The implemented circuit contained self-activation and lateral inhibition, with two diffusible quorum-sensing signals [16]. Second, solitary patterns were engineered in HEK cells using the Nodal–Lefty system [17]. The issue is that small molecules have roughly the same diffusion constant, although extra nonspecific binding can help slow down diffusion. In recent work, a more robust three-node network was implemented using synthetic circuits of six genes with small diffusible quorum-sensing molecules and extra control cassettes [18]. This showed regular patterns in growing bacterial colonies, but the variability is immense, and robustness is still limited. Hence, while there is progress in the rational design of circuits with specific properties, the lack of control over robustness represents a significant downside in tissue engineering, patterned biomaterial deposition, and bridging developmental programs [19][20][21].
@@
The need to develop theoretical frameworks beyond the two-equation Turing model has been recognized previously. An exhaustive exploration of all two- and three-node topologies showed that three nodes are, on average, more robust than two nodes, pointing toward complexity increasing robustness [22]. More specifically, approximately 60% of all topologies produced patterns for some parameter combinations, but the parameter space producing Turing patterns was overall minuscule, around 0.1%. This finding is supported by another study that explored networks with up to four nodes [23]. Motivated by Robert May’s work in theoretical ecology [24] (based on earlier work by Eugene Wigner [25]) a random matrix approach for $N\leq 6$ showed that larger networks add robustness by reducing the ”diffusive threshold”, i.e., softening the requirement of differential diffusion [26]. This raises the question: why not investigate even larger matrices? While the diagonalization of Jacobian matrices is a slow $O(N^{3})$ process, it is still efficient for significantly larger networks with modern computers. The advantages of exploring random Jacobian matrices are apparent; such an approach produces excellent statistics, and biological realism can be introduced through the distributions and topologies.
@@
Here, we go beyond small Turing networks and use large random Jacobian matrices to systematically explore how network size affects robustness. We begin by motivating distributions for Jacobian matrix elements from explicit small Hill-function-based models for gene regulation, as relevant to synthetic biology [18]. We then continue by exploring large random networks with up to $N=100$ nodes using corresponding random Jacobian matrices with two diffusers and variable sparsity. We identify an optimal network size $N_{\text{opt}}\sim 5-8$, arising from a tradeoff between stability without diffusion and instability with diffusion, each with opposite $N$-dependencies. This increased robustness relieves the constraints on parameters, known as Turing conditions, including differential diffusion for the two diffusing species. Our results are expected to renew the quest for robust Turing patterns in synthetic biology and to identify Turing modules in larger networks of developmental biology.
@@
Figure 1: Network graph representations of pre-defined Turing topologies. (Ai) 2-node Gierer-Meinhardt network. (Bi) 3-node network based on [22, 18]. (C) Extended 4-node network. (D) Large $N>\!\!>1$ network. (E) Number of different Turing topologies (graphs) for $N$ nodes, exemplified by all 2 and 3-node networks in (Aii) and (Bii), respectively. Dark blue (red) arrows indicate activation (inhibition), and wiggly lines indicate diffusion.
@@ Results
We are interested in the robustness of large Turing networks, that is reaction-diffusion models with many molecular species with a large parameter space to support diffusion-driven instabilities. Fig. 1 shows example networks, ranging from small ($N=2$ nodes) to large ($N>\!\!>1$). We restrict ourselves to two diffusing species similar to Turing’s original paper. In order to understand how to model large networks we begin by investigating examples of smaller networks to gain intuition. For a network of $N$ nodes and hence $N$ species with time and space dependent concentrations $\bm{X}(\bm{r},t)=\{x_{1}(\bm{r},t),x_{2}(\bm{r},t),\dots,x_{N}(\bm{r},t)\}$, the general dynamics are given by the partial differential equations
$\frac{\partial x_{i}}{\partial t}=f_{i}(\bm{X};\bm{\theta})+D_{i}\nabla^{2}\bm{X},$ (1)
where $\bm{D}=diag(D_{1},D_{2},0,\dots,0)$ is the diagonal diffusion matrix, and Laplacian $\nabla^{2}=\sum_{i=1}^{d}\partial^{2}/\partial r_{i}^{2}$ in $d$ dimensions, given by the sum of second spatial derivatives. For simplicity, we assume an infinite domain, allowing us to neglect boundary effects and to use a continuous wave number. As our motivation is to learn how to build these with synthetic gene circuits, we model the activating and inhibiting interactions with Hill functions, along with basal expression and degradation:
@@
To get rid off the second derivatives, we Fourier transform in the space domain, or more simply apply a wave-like perturbation $\delta x(\bm{r},t)=\delta\tilde{x}(t)\exp{(i\bm{k}\bm{r})}$ with $\bm{k}=\{k_{1},k_{2},\dots,k_{d})$ and $\bm{k}\bm{r}$ the scalar product between the $\bm{k}$ and $\bm{r}$ vectors. Plugged into Eq. 4, this leads to a new Jacobian matrix with modified diagonal matrix elements and an extra parameter $k=|\bm{k}|$
@@
indicating that for a rotationally invariant system, the dependence is only on the modulus of $\bm{k}$, rendering linear stability analysis an effective one-dimensional problem. As a result, diagonalization of the modified Jacobian matrix leads to a $k$-dependent eigenvalues, called dispersion relations, where the one with the largest real part, termed $Re(\lambda_{\text{max}}(k))$, determines the stability. For a Turing instability, we require the homogeneous steady state without diffusion to be stable (corresponding to $k$=0), i.e. with a negative real part given by $Re(\lambda_{\text{max}}(k(0)))<0$. Furthermore, with diffusion we require instability for a wave number ($k>0$) and thus a positive real part given by $Re(\lambda_{\text{max}}(k))>0$. For a classic Turing instability leading to a well-defined wave pattern, there is a $k_{\text{max}}>0$, for which the dispersion has a clear maximum, and for $k\rightarrow\infty$, the system becomes stable again (negative real part of eigenvalue). This type of instability is called Turing I (Fig. 2A). However, there are other possibilities, such as instability for $k\rightarrow\infty$ (Turing II in Fig. 2B). (Note, this is a simplified classification scheme - in other schemes our Turing I is given by Turing Ia, and our Turing II is given by Turing Ib, IIa, and IIb with subtle differences [22].) Yet, another type of instability is the Turing-Hopf, for which the imaginary part can additionally lead to oscillations (not shown).
@@
Focusing on small networks ($N=2-4$), we used Latin hypercube sampling of $10^{7}$ parameter sets and filtered these for Turing I instabilities. Fig. 3 shows the fits of the empirical histograms of the Jacobian matrix elements ($j$) to the beta distribution $B(j;\alpha,\beta)$ (see Methods for details on parameter searches and fitting procedure). The beta distribution depends on two parameters, $\alpha$ and $\beta$, allowing a wide range of distributions from symmetric Gaussian-like to non-Gaussian shapes with various levels of skewness and kurtosis. We generally observe that the Turing-I matrix elements are more narrowly distributed in line with Turing I having to fulfill Turing conditions, leading to a subset of parameters and hence Jacobian matrix elements. This can clearly be seen in Fig. 3A (see Fig. S1 in Supplementary Materials for additional histograms). To fulfill the Turing conditions, the (1,1) Jacobian matrix elements need to be positive (activator), which is the case for the histogram and fitted line (blue), but not for the orange line, representating the non-Turing case. Furthermore, the (2,2) matrix elements need to be negative (inhibitor), which is here already fulfilled for non-Turing due to model constraints (degradation). In contrast, the (1,2) and (2,1) matrix elements can respectively be either negative and positive (here the case even for non-Turing due to model constraints), or the other way around (however not for our 2-equation model) [6]. We generally did not observe multimodal distributions, pointing to a certain level of simplicity in the distributions.
@@
The difference between Turing and non-Turing Jacobian matrix elements is further visible in Fig. 4, where panel A shows beehive plots based on the empirical data from Fig. 3. Clearly, the variances of the off-diagonal matrix elements of the non-Turing cases are much broader than the Turing-I cases. This is further provided in panel B, showing the sum of the Jacobian matrix elements (equivalent plots can be made for the mean of those random variables). For larger matrices, we expect to observe the central limit theorem, with the sums calculated from $N(N-1)$ off-diagonal Jacobian matrix elements. For increasing matrix size, and this is already evident for $N=4$, we obtain approximate Gaussian distributions (still with a sharp peak at zero): The mean is approximately zero, the variance approaches a finite value (as variances are positive numbers), and the skewness and kurtosis vanish. Hence not surprisingly, on average a Gaussian distribution with zero mean and finite variance describes the Jacobian matrix elements. Note this trend would be even stronger, if we average over all possible network topologies with $N$ nodes, which scales as $N_{G}=3^{N}9^{\binom{N}{2}}$ and hence is super-exponentially increasing with $N$ (see Fig. 1E). In the following we exploit these insights and use random Jacobian matrices directly.
@@
Figure 4: Jacobian matrix elements are different for Turing I and non-Turing. (A) Beehive plots of the mean (top) and variance (bottom) of the empirical distributions of the non-zero off-diagonal Jacobian matrix elements from Fig. 3 (based on the pre-defined 2-4-node network topologies in Fig. 1Ai, Bi, and C). A statistical Brown-Forsythe test was utilised to verify whether the difference in variance between each pair of Turing and non-Turing distributions was statistically significant. For all pairs of the Jacobian distributions except one, we found that the test showed a statistically significant difference between the variances. Furthermore, 12 out of 16 $p$-values were under 0.001, showing a strong statistical significance. (B) Sum of off-diagnal matrix elements, showing a clear visual difference between Turing I (light blue) and non-Turing (light orange) for the 2-4 node topologies. The distributions become increasingly symmetric (and Gaussian) for more nodes.
Inspired by Robert May‘s statistical treatment of non-equilibrium ecological communities that studied asymmetric random Jacobians to represent linearized population dynamics, as opposed to Wigner’s symmetric (Hermitian) random matrices [24][27], we analyzed Turing instabilities by sampling asymmetric random Jacobian matrices that represent the linearized reaction dynamics. Similar to May’s ”neutral interaction” model, our Jacobian matrix without diffusion is given by
@@
where $\bm{G}$ is a random matrix with elements $g_{ij}$ randomly assigned from a Gaussian distribution that has mean value zero and variance Var$(g_{ij})=\sigma^{2}$, but with diagonal terms $g_{ii}=0$. For simplicity (and convenience), the latter are modeled by the identity matrix, $\bm{I}$, responsible for providing stability. In such a Turing system, a positive value of $J_{ij}$ (or $g_{ii}$) implies activation while a negative value implies inhibition, similarly to previous matrix representations of Turing systems [22][28]. The identity matrix $\bm{I}$ represents the degradation of the species. Further self-interactions are neglected, leading to the curious result that there are no Turing instabilities for $N=2$ (see proof in Methods). While the Gaussian distribution is motivated by Fig. 4, there is no reason to expect the reaction kinetics in a reaction-diffusion system to be independent as there is no reason for the population dynamics in the work of May [24] to be independent (or normally distributed). Nevertheless, in the absence of experimental understanding of what the distribution of the parameters should be, the potential of the random matrix approach to explaining stability principles is greatly evidenced for population dynamics [29, 30] and in small Turing systems ($N\leq 6$) [26].
@@
Next, we investigated the eigenvalue distributions of such random matrices (see Fig. 5A). In the absence of diffusion (i.e. at $k=0$), the eigenvalues are distributed in a circle with radius $\gamma=\sigma\sqrt{N}$ according to May’s ”circular law”, derived from random ecological communities [24][27]. In the presence of diffusion, the eigenvalue distribution can be described as having ”in-bulk” and ”outlier” distributions (see Fig. 5B). The outlier distribution has extreme negative real parts, following approximately $-N-k^{2}\sum_{i=1}^{N}D_{i}$ (see Methods), which tends to minus infinity for increasing $k$ and network size $N$. As we are only interested in the maximal real part to determine the linear stability, it is safe to ignore the ”outliers” and only focus on the “in-bulk” eigenvalue distribution where the maximal real eigenvalues are located in.
@@
Figure 5: Eigenvalue distribution of a large random network. (A) Circular scatter plot and corresponding kernel density estimations of eigenvalues in the complex plain for network size $N=100$, variance $\sigma^{2}=0.01$, and diffusion constants $D_{1}=1$, $D_{2}=10$. While the system without diffusion has all the eigenvalues distribute around $\gamma=1$, the same system with diffusion has most but not all of its eigenvalues distribute in a circle with radius $\gamma=1$. (B) Full eigenvalue distribution with outliers. Green, red and blue points correspond to wave numbers $k<1$, $k=1$ and $k>1$, respectively.
@@
We subsequently sampled thousands of random matrices to analyse how network size affects the stability of a reaction-diffusion system. We observed three relationships between network size and stability. Firstly, in the absence of diffusion, a random reaction-diffusion system with fixed radius $\gamma$ is likely more stable for smaller network size (Fig. 6A). Asymptotically, the circlar law predicts a step function for $N\rightarrow\infty$, i.e. stability for radii $\gamma<1$ (as circles are fully to the left of the imaginary axis) and instability for $\gamma>1$ (as circles crosses the imaginary axis) [25]. Secondly, a previously stable system without diffusion is more likely to turn unstable (i.e. have a diffusion-driven instability) when diffusion is present for larger network size (Fig. 6B). Therefore, thirdly, out of all random matrices, Turing instabilities are maximized (i.e. found the most) at an optimal intermediate network size that is not too small or not too large (Fig. 6C).
@@
As the Turing-Hopf instability is competing with the Turing instability, and Turing I is of importance in producing spatial patterns, we analysed the effects of network size on Turing I and Turing-Hopf instabilities. Out of stable systems without diffusion, the systems are likely to become Turing-Hopf unstable in the presence of diffusion with larger network size (Fig. 6B). This trend for the Turing-Hopf instability is similar to the trend for Turing I. However, particularly at a very large network size, the Turing-Hopf instability becomes more likely to occur and ultimately out-competes Turing I, as can be seen for $N=50$ in the figure. A similar trend is observed for Turing I relative to the overall Turing instabilities of all types: Turing I is more likely for larger network size being approximately half the occurrence of all Turing instabilities. Among all Turing instability types, we found that Turing II is the most common to arise (4.2% maximum occurrence of all random reaction-diffusion systems), followed by Turing I (3.6%).

### v2

@@ Abstract
Many cellular patterns exhibit a reaction-diffusion component, suggesting that Turing instability may contribute to pattern formation. However, biological gene-regulatory pathways are more complex than simple Turing activator-inhibitor models and generally do not require fine-tuning of parameters as dictated by the Turing conditions. To address these issues, we employ random matrix theory to analyze the Jacobian matrices of larger networks with robust statistical properties. Our analysis reveals that Turing patterns are more likely to occur by chance than previously thought and that the most robust Turing networks have an optimal size, consisting of only a handful of molecular species, thus significantly increasing their identifiability in biological systems. Broadly speaking, this optimal size emerges from a trade-off between the highest stability in small networks and the greatest instability with diffusion in large networks. Furthermore, we find that with multiple immobile nodes, differential diffusion ceases to be important for Turing patterns. Our findings may inform future synthetic biology approaches and provide insights into bridging the gap to complex developmental pathways.
@@ Introduction
Spatial patterns and structures are common in biological systems, ranging from microbial communities and biofilms to developmental biology and ecological systems [1] [2] [3]. In a seminal work, Alan Turing proposed a self-organizing, emergent mechanism for pattern formation based on a diffusion-driven instability [4], later formalized by Gierer and Meinhardt in terms of a slowly diffusing activator and a fast-diffusing inhibitor molecule [5]. According to Turing’s definition, the chemical reactions are stable without diffusion, leading to a homogeneous steady state, but become unstable with diffusion, forming a periodic pattern at certain wave numbers. The mathematical analysis is generally based on reaction-diffusion models, implemented using partial differential equations, and linear stability analysis to investigate the effect of perturbations on a steady state [6]. Nonetheless, employing Turing models to understand biological patterns is still a topic of debate.
@@
The Turing mechanism clearly has biological relevance in cell and developmental biology, likely explaining aspects of digit formation in mice, scale-patterns in zebrafish, fingerprint formation, and cortical folds of the fetal brain [7][8][9][10]. However, this mechanism has two main drawbacks: extreme simplicity and a need for fine-tuning. The latter requires parameters to be carefully chosen to produce patterns. In contrast to simple activator-inhibitor Turing models, biological gene regulatory networks in embryonic development generally consist of hundreds of molecular species and are notable for their immense complexity, hierarchical structure, and tolerance to noise [11]. In contrast, the Turing conditions lead to a lack of structural robustness, which is at odds with the noise tolerance and evolutionary adaptability required for such patterning solutions to occur [12][13][14]. However, there are also bottom-up approaches to further elucidate the issues with Turing models.
@@
As developmental biology is complex, synthetic biology provides an alternative route, following Feynman’s mantra: “What I cannot create, I do not understand” [15]. By implementing the Turing mechanism from scratch in cells that communicate via chemicals, our understanding of how to generate stable pattern can be systematically improved. Two previous synthetic biology attempts failed due to the lack of differential diffusion, leading to irregular patterns. First, stochastic Turing patterns were engineered in E. coli cells. The implemented circuit contained self-activation and lateral inhibition, with two diffusible quorum-sensing signals [16]. Second, solitary patterns were engineered in HEK cells using the Nodal–Lefty system [17]. The issue is that small molecules have roughly the same diffusion constant, although extra nonspecific binding can help slow down diffusion. In a recent work, a more robust three-node network was implemented using synthetic circuits of six genes with small diffusible quorum-sensing molecules and extra control cassettes [18]. This showed regular patterns in growing bacterial colonies, but the variability is immense, and robustness is still limited. Methods to increase the reproducibility of these patterns, such as solving the inverse problem  given an experimental pattern, have also been recently developed[19]. Hence, while there is progress in the rational design of circuits with specific properties, the lack of control over robustness represents a significant downside in tissue engineering, patterned biomaterial deposition, and bridging developmental programs [20][21][22].
@@
The need to develop theoretical frameworks beyond the two-equation Turing model has been recognized previously. An exhaustive exploration of all two- and three-node topologies showed that three nodes are, on average, more robust than two nodes, pointing toward complexity increasing robustness [23]. Note that ”nodes” refers here to genes and links to regulatory interactions. These networks should not be confused with Turing models on graphs, which deal with discrete topologies, not continuous space [24] [25] [26] [27]. More specifically, approximately 60% of all topologies produced patterns for some parameter combinations, but the parameter space producing Turing patterns was overall minuscule, around 0.1%. This finding is supported by another study that explored networks with up to four nodes [28]. Motivated by Robert May’s work in theoretical ecology [29] (based on earlier work by Eugene Wigner [30]) a random matrix approach for $N\leq 6$ showed that larger networks add robustness by reducing the ”diffusive threshold”, i.e., softening the requirement of differential diffusion [31]. This raises the question: why not investigate even larger matrices? While the diagonalization of Jacobian matrices is a slow $O(N^{3})$ process, it is still efficient for significantly larger networks with modern computers. The advantages of exploring random Jacobian matrices are apparent; such an approach produces excellent statistics, and biological realism can be introduced through the distributions and topologies.
@@
Here, we go beyond small Turing networks and use large random Jacobian matrices to systematically explore how network size affects robustness of patterning continuous space. We begin by motivating distributions for Jacobian matrix elements from explicit small Hill-function-based models for gene regulation, as relevant to synthetic biology [18]. We then continue by exploring large random networks with up to $N=100$ nodes using corresponding random Jacobian matrices with two diffusers and variable sparsity. We identify an optimal network size $N_{\text{opt}}\sim 5-8$, arising from a tradeoff between stability without diffusion and instability with diffusion, each with opposite $N$-dependencies. This increased robustness relieves the constraints on parameters, known as Turing conditions, including differential diffusion for the two diffusing species. Our results are expected to renew the quest for robust Turing patterns in synthetic biology and to identify Turing modules in larger networks of developmental biology.
@@
Figure 1: Network graph representations of pre-defined Turing topologies. (Ai) 2-node Gierer-Meinhardt network. (Bi) 3-node network based on [23, 18]. (C) Extended 4-node network. (D) Large $N>\!\!>1$ network. (E) Number of different Turing-network topologies for $N$ nodes, exemplified by all 2 and 3-node networks in (Aii) and (Bii), respectively. Dark blue (red) arrows indicate activation (inhibition), dashed lines denote missing links, and wiggly lines indicate diffusion.
@@ Results
We are interested in the robustness of large Turing networks, that is reaction-diffusion models with many molecular species with a large parameter space to support diffusion-driven instabilities. Fig. 1 shows example networks, ranging from small ($N=2$ nodes) to large ($N>\!\!>1$). We restrict ourselves to two diffusing species similar to Turing’s original paper. In order to understand how to model large networks we begin by investigating examples of smaller networks to gain intuition. For a network of $N$ nodes and hence $N$ species with time and space dependent concentrations $\bm{X}(\bm{r},t)=\{x_{1}(\bm{r},t),x_{2}(\bm{r},t),\dots,x_{N}(\bm{r},t)\}$, the general dynamics are given by the following partial differential equations
$\frac{\partial x_{i}}{\partial t}=f_{i}(\bm{X};\bm{\theta})+D_{i}\nabla^{2}x_{i},$ (1)
where $\bm{D}=diag(D_{1},D_{2},0,\dots,0)$ is the diagonal diffusion matrix, and Laplacian $\nabla^{2}=\sum_{i=1}^{d}\partial^{2}/\partial r_{i}^{2}$ in $d$ dimensions, given by the sum of second spatial derivatives. For simplicity, we assume an infinite domain, allowing us to neglect boundary effects and to use a continuous wave number. The nonlinearities are given by the interactions between the nodes. As shown in Fig. 1Aii and Bii, we consider all possibilities for each connection, depicted as black for activation, red for inhibition and dotted for no interaction. We will refer to activation and inhibition nodes as positive and negative edges respectively. Curved lines denote diffusion, which as mentioned will only be considered for the first two nodes. As our motivation is to learn how to build these with synthetic gene circuits, we model the activating and inhibiting interactions with Hill functions, together with basal expression and degradation terms:
@@
To remove the second derivatives, we Fourier transform in the space domain, or more simply apply a wave-like perturbation $\delta x(\bm{r},t)=\delta\tilde{x}(t)\exp{(i\bm{k}\bm{r})}$ with $\bm{k}=\{k_{1},k_{2},\dots,k_{d})$ and $\bm{k}\bm{r}$ the scalar product between the $\bm{k}$ and $\bm{r}$ vectors. Plugged into Eq. 4, this leads to a new Jacobian matrix with modified diagonal matrix elements and extra parameter $k=|\bm{k}|$, given by
@@
indicating that for a rotationally invariant system, the dependence is only on the modulus of $\bm{k}$, rendering linear stability analysis an effective one-dimensional problem. As a result, diagonalization of the modified Jacobian matrix leads to $k$-dependent eigenvalues, called dispersion relations, where the one with the largest real part, termed $Re(\lambda_{\text{max}}(k))$, determines the stability. For a Turing instability, we require the homogeneous steady state without diffusion to be stable (corresponding to $k$=0), i.e. with a negative real part given by $Re(\lambda_{\text{max}}(k(0)))<0$. Furthermore, with diffusion, we require instability for a wave number ($k>0$) and thus a positive real part given by $Re(\lambda_{\text{max}}(k))>0$. For a classic Turing instability leading to a well-defined wave pattern, there is a $k_{\text{max}}>0$, for which the dispersion has a clear maximum, and for $k\rightarrow\infty$, the system becomes stable again (negative real part of eigenvalue). This type of instability is called Turing I (Fig. 2A). However, there are other possibilities, such as the instability for $k\rightarrow\infty$ (Turing II in Fig. 2B). In the latter case, there is no well defined wavelength for pattern formation, and hence Turing II cases are generally not considered in the following. (Note that this is a simplified classification scheme - in other schemes our Turing I is given by Turing Ia, and our Turing II is given by Turing Ib, IIa, and IIb with subtle differences [23].) Yet, another type of instability is the Turing-Hopf, for which the imaginary part can additionally lead to oscillations (not shown).
@@
Focusing initially on small networks ($N=2-4$), we employed Latin hypercube sampling of $10^{7}$ parameter sets and filtered them for Turing I instabilities. Fig. 3 shows the fits of the empirical histograms of the Jacobian matrix elements ($j$) to the beta distribution $B(j;\alpha,\beta)$ (see Methods for details on parameter searches and fitting procedure). The beta distribution depends on two parameters, $\alpha$ and $\beta$, allowing a wide range of distributions from Gaussian-like symmetric to non-Gaussian shapes with various levels of skewness and kurtosis. We generally observe that the Turing-I matrix elements are more narrowly distributed in line with Turing I having to fulfill Turing conditions, leading to a subset of parameters and hence Jacobian matrix elements. This can be clearly seen in Fig. 3A (see Fig. S1 in the Supplementary Figures for additional histograms). To fulfill the Turing conditions, the (1,1) Jacobian matrix elements need to be positive for activation, which is the case for the histogram and fitted line (blue), but not for the orange line, representing the non-Turing case. Note, while Node 1 is an activator, degradation contributes to the diagonal matrix elements, potentially leading to negative values. Furthermore, the (2,2) matrix elements need to be negative for inhibition, which is here already fulfilled for non-Turing due to model constraints (degradation). In contrast, the (1,2) and (2,1) matrix elements can, respectively, be either negative and positive (here the case even for non-Turing due to model constraints), or the other way around (however not for our 2-equation model) [6]. We generally did not observe multimodal distributions, which indicates a certain level of simplicity in the distributions.
@@
The difference between Turing and non-Turing Jacobian matrix elements is further visible in Fig. 4, where panel A shows beehive plots based on the empirical data from Fig. 3. Clearly, the variances of the off-diagonal matrix elements of the non-Turing cases are much broader than the Turing-I cases. This is further illustrated in panel B, showing the sum of the Jacobian matrix elements (equivalent plots can be made for the mean of those random variables). For larger matrices, we expect to observe the central limit theorem, with the sums calculated from $N(N-1)$ off-diagonal Jacobian matrix elements. For increasing matrix size, this trend is already evident for $N=4$, although there is still a significant cusp at zero: The mean is approximately zero, the variance approaches a finite value (as variances are positive numbers), and the skewness and kurtosis vanish. Hence, not surprisingly, on average a Gaussian distribution with zero mean and finite variance describes the sum of Jacobian matrix elements. Note that this trend would be even stronger, if we average over all possible network topologies with $N$ nodes, which scales as $N_{G}=3^{N}9^{\binom{N}{2}}$ and hence is superexponentially increasing with $N$ (see Fig. 1E). In the following, we exploit these insights and use random Jacobian matrices directly.
@@
Figure 4: Jacobian matrix elements are different for Turing I and non-Turing. (A) Beehive plots of the mean (top) and variance (bottom) of the empirical distributions of the non-zero off-diagonal Jacobian matrix elements from Fig. 3 (based on the pre-defined 2-4-node network topologies in Fig. 1Ai, Bi, and C). A statistical Brown-Forsythe test was utilized to verify whether the difference in variance between each pair of Turing and non-Turing distributions was statistically significant. For all pairs of the Jacobian distributions except one, we found that the test showed a statistically significant difference between the variances. Furthermore, 12 out of 16 $p$-values were under 0.001, showing a strong statistical significance. (B) Sum of off-diagnal matrix elements, showing a clear visual difference between Turing I (light blue) and non-Turing (light orange) for the 2-4 node topologies. The distributions become increasingly symmetric, and ultimately Gaussian, for large numbers of nodes.
Inspired by Robert May‘s statistical treatment of non-equilibrium ecological communities that studied asymmetric random Jacobians to represent linearized population dynamics, contrasting Wigner’s symmetric (Hermitian) random matrices [29][32], we analyzed Turing instabilities by sampling asymmetric random Jacobian matrices that represent the linearized reaction dynamics. Similar to May’s ”neutral interaction” model, our Jacobian matrix without diffusion is given by
@@
where $\bm{G}$ is a random matrix with elements $g_{ij}$ randomly assigned from a Gaussian distribution that has mean value zero and variance Var$(g_{ij})=\sigma^{2}$, but with diagonal terms $g_{ii}=0$. For simplicity (and convenience), the latter are modeled by the identity matrix, $\bm{I}$, responsible for providing stability. In such a Turing system, a positive value of $J_{ij}$ (or $g_{ii}$) implies activation while a negative value implies inhibition, similarly to previous matrix representations of Turing systems [23][33]. The identity matrix $\bm{I}$ represents the degradation of the species. Further self-interactions are neglected, leading to the curious result that there are no Turing instabilities for $N=2$ (see proof in Methods). While the Gaussian distribution is motivated by Fig. 4, there is no reason to expect the reaction kinetics in a reaction-diffusion system to be independent as there was no reason for the population dynamics in the work of May [29] to be independent. However, detailed understanding of realistic models and their parameters, or Jacobian matrices of linearized models is missing. Hence, the statistical random-matrix approach has been powerful in understanding the principles of stability in population dynamics [34, 35] and small Turing systems ($N\leq 6$) [31].
@@
Next, we investigate the eigenvalue distributions of such random matrices (see Fig. 5A). In the absence of diffusion (i.e. at $k=0$), the eigenvalues are distributed in a circle with radius $\gamma=\sigma\sqrt{N}$ according to May’s ”circular law”, derived from random ecological communities [29][32] and exact in the asymptotic limit. In the presence of diffusion, the eigenvalue distribution can be described as having ”in-bulk” and ”outlier” distributions (see Fig. 5B). The outlier distribution has extreme negative real parts, following approximately $-N-k^{2}\sum_{i=1}^{N}D_{i}$ (see Methods), which tends to minus infinity for increasing $k$ and network size $N$. As we are only interested in the maximal real part to determine the linear stability, it is safe to ignore the ”outliers” and only focus on the “in-bulk” eigenvalue distribution, containing the maximal real eigenvalues.
@@
Figure 5: Eigenvalue distribution of a large random network. (A) Circular scatter plot and corresponding kernel density estimations of eigenvalues in the complex plain (green at the top and blue on the right) for network size $N=100$, variance $\sigma^{2}=0.01$, and diffusion constants $D_{1}=1$, $D_{2}=10$. $k$-dependent movement of the eigenvalues is visible, and finite-$N$ effects in the kernel density estimates as well, such as suppression of the density at the boundary of the circle. While the system without diffusion has all the eigenvalues distribute around $\gamma=1$, the same system with diffusion has most but not all of its eigenvalues distribute in a circle with radius $\gamma=1$. (B) Full eigenvalue distribution at a large scale with focus on the outliers (the eigenvalue circle now appears like a vertical line.) Green, red and blue points correspond to wave numbers $k<1$, $k=1$ and $k>1$, respectively.
@@
We subsequently sampled thousands of random matrices to analyse how network size affects the stability of a reaction-diffusion system. We observed three relationships between network size and stability. Firstly, in the absence of diffusion, a random reaction-diffusion system with fixed radius $\gamma$ is likely more stable for smaller network size (Fig. 6A). Asymptotically, the circular law predicts a step function for $N\rightarrow\infty$, i.e. stability for radii $\gamma<1$ (as circles are fully to the left of the imaginary axis) and instability for $\gamma>1$ (as circles cross the imaginary axis) [30]. Secondly, a previously stable system without diffusion is more likely to turn unstable (i.e. having a diffusion-driven instability) when diffusion is present for larger network size (Fig. 6B). Therefore, thirdly, out of all random matrices, Turing instabilities are maximized (i.e. found the most) at an optimal intermediate network size (Fig. 6C).
@@
As the Turing-Hopf instability is competing with the Turing instability, and Turing I is of importance in producing spatial patterns, we analyzed the effects of network size on Turing I and Turing-Hopf instabilities. Out of stable systems without diffusion, the systems are likely to become Turing-Hopf unstable in the presence of diffusion with larger network size (Fig. 6B). This trend for the Turing-Hopf instability is similar to the trend for Turing I. However, particularly at a very large network size, the Turing-Hopf instability becomes more likely to occur and ultimately out-competes Turing I, as can be seen for $N=50$ in the figure. A similar trend is observed for Turing I relative to the overall Turing instabilities of all types: Turing I is more likely for larger network size being approximately half the occurrence of all Turing instabilities. Among all Turing instability types, we found that Turing II is the most common to arise (4.2% maximum occurrence of all random reaction-diffusion systems), followed by Turing I (3.6%).